OpenStudy (nanii215):

f(x)= 1/x-2, g(x)=2x+1/x A. Use composition to prove whether or not the functions are inverses of each other. B. Express the domain of the compositions using interval notation.

OpenStudy (caseysorensen):

i think it might be B.Express the domain of the compositions using interval notation.

OpenStudy (nanii215):

a & b is what i need to do.

OpenStudy (caseysorensen):

oh ok

OpenStudy (caseysorensen):

A. combination at the point x=4. f(4)=5(4)+2=22 and g(4)=42-1=15 ... (f·g)(x), (5x+2)*(x2-1) =5x3+2x2-5x-2, (f·g)(4), 5(43)+2(42)-5(4)-2 .B f (x) = 2x and g(x) = x are inverse functions because f o g(x) = f (g(x)) = f ( x) = 2( x) = x and g o f (x) = g(f (x)) = g(2x) = (2x) = x . Similarly, f (x) = x + 1